OpenStudy (anonymous):

solve each literal equation for the given variable. You must show your work. Solve S = 4v2 for v. Solve M = 2x + 3y for y.

OpenStudy (anonymous):

is the first one s=4v^2

OpenStudy (anonymous):

Is the 2 an exponent?

OpenStudy (anonymous):

yes, i believe so

OpenStudy (anonymous):

Alright, you wanna undo operations on the right so you can get v alone. to get rid of the 4 you divide by 4 on both sides s/4=v^2 to eliminate the square you take the square root of both sides \[\pm \sqrt{S/4}=v\]

OpenStudy (anonymous):

ok(: